A company selects five samples (kequals=5​) of size four (nequals=4​) parts used in a production process and measures the diameter of each part. Develop an

x overbarx chart to determine if the process is stable or not

Sample

Obs1

Obs2

Obs3

Obs4

1

6

7

4

7

2

5

6

7

7

3

9

8

7

7

4

6

7

7

6

5

4 4    6    5

n

Upper A 2A2

2

1.880

3

1.023

4

0.729

5

0.577

6

0.483

The upper control limit is

?
Answer #1

Solution:

For X_bar Chart the UCL is = X-bar_bar + A2 * R_bar

R_bar is the mean of ranges of all the 5 levels and A2 is a constant which depends on the sample size.

Hence,

R_bar = (3+2+2+1+2)/5 = 2

A2 = 0.729 (constant for n = 4)

X_bar_bar = ( 6 + 6.25+.7.75+6.5+4.75)/5 = 31.25/5 = 6.25

therefore,

UCL = X-bar_bar + A2 * R_bar = 6.25 + 0.729 * 2 = 6.25 + 1.458 = 7.708